OpenStudy (anonymous):

I need my work checked: If you have the equation: x^2+6x-2=0 Would x equal -7, 1

OpenStudy (watchmath):

no

OpenStudy (anonymous):

How would I solve it then?

OpenStudy (watchmath):

Quadratic formula or completing the square.

OpenStudy (anonymous):

That is what I have been using. This is my work so far. x\[x ^{2 } +6x -2+0\] \[-2=2\]

OpenStudy (anonymous):

\[x ^{2} +6x=2 6x/2 =3 3 squared equals 9 x squared plus 6x+9=2+9 x squared plus 6x plus 9=11 (x+3)quantity squared equals 11 x+3 equals square root of 11 x+3=+ minus 3.3 x=(-3) minus 4 x=(-3)+4 (-3)-4=(-)7 (-)3+4=1

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}\div 2a\] That's the quadratic formula, you would use it to find the x-value of each x-intercept. (You would do the division last, it's hard to note this in the equation input thing) \[[-6\pm \sqrt{6^{2}-4\left( 1 \right)\left( -2 \right)}] \div2(1)\]\[-6\pm \sqrt{36+8} \div 2\]\[-6 \pm \sqrt{44} \div 2\] Now because 44 is not a perfect square number, both of your answers will be real numbers but not whole numbers. \[(-6 + 6.63) \div 2\]\[.63 \div2 \approx .32\] and the other x-value:\[(-6 - 6.63) \div 2\]\[-12.63 \div2 =-6.32\]

OpenStudy (anonymous):

You were up until: x+3=+ minus 3.3 then you got a 4 in there for no reason. x1=-3-3.3...=-6.3... and x2=-3+3.3...=+0.3... which you can see iris31 also got using the quadratic formula.

OpenStudy (anonymous):

Right, and if you have an iPhone or something like that there's a great graphing calculator app called Free Graphing Calculator. It helps immensely when you want to check your answers (I used it to check mine; it helped me catch a mistake I made)